# Generalized Bayesian Inference for Causal Effects using Covariate Balancing Procedure

### v2

@@ Appendix A Simulation Results
Confounders Logit 0.093 6.197 0.933 0.231 0.441 1.005 0.221 3.841 0.947 0.148 0.786 1.002 0.053 2.699 0.950 0.105 2.650 1.000
CBPS 0.240 6.207 0.931 0.230 1.137 1.006 0.296 3.846 0.945 0.148 1.053 1.003 0.014 2.703 0.949 0.105 0.700 1.001
RCAL CV 0.337 5.901 0.920 0.208 1.597 0.957 0.066 3.768 0.919 0.132 0.235 0.983 0.112 2.676 0.920 0.093 5.600 0.991
$\lambda=0.01$ 0.340 6.118 0.912 0.208 0.027 3.713 0.924 0.132 0.176 2.601 0.928 0.093
$\lambda=0.05$ 0.690 5.431 0.942 0.208 0.945 3.367 0.947 0.131 1.263 2.362 0.952 0.093
$\lambda=0.10$ 0.654 5.175 0.950 0.205 0.759 3.236 0.952 0.129 0.972 2.267 0.957 0.091
BRCAL PCIC 0.211 6.168 0.932 0.229 Ref. Ref. 0.281 3.835 0.941 0.146 Ref. Ref. 0.020 2.700 0.946 0.105 Ref. Ref.
$\omega=0.2$ 0.001 5.932 1.000 0.621 0.149 3.777 1.000 0.392 0.086 2.680 1.000 0.277
$\omega=0.5$ 0.129 6.096 0.999 0.393 0.246 3.818 0.999 0.248 0.040 2.694 0.999 0.175
$\omega=1.0$ 0.182 6.149 0.976 0.278 0.281 3.831 0.978 0.175 0.026 2.699 0.977 0.124
$\omega=1.5$ 0.212 6.169 0.930 0.227 0.280 3.836 0.938 0.143 0.020 2.700 0.939 0.101
Predictors Logit 0.024 11.122 0.872 0.289 0.329 1.068 0.002 7.901 0.912 0.226 0.007 1.109 0.462 6.088 0.932 0.183 2.232 1.146
CBPS 0.030 10.645 0.411 1.022 0.318 7.193 0.914 0.246 1.178 1.010 0.230 5.339 0.897 0.174 1.111 1.005
RCAL CV 1.465 6.231 0.916 0.209 20.068 0.598 1.321 5.575 0.827 0.132 4.893 0.782 1.078 4.920 5.208 0.926
$\lambda=0.01$ 5.587 20.343 0.679 23.886 3.638 12.275 0.600 0.152 1.817 7.333 0.610 0.093
$\lambda=0.05$ 1.848 9.896 0.817 0.218 1.040 4.919 0.856 0.133 0.829 3.141 0.875 0.095
$\lambda=0.10$ 1.500 5.839 0.929 0.209 1.859 3.613 0.931 0.132 2.126 2.490 0.938 0.093
BRCAL PCIC 0.073 10.414 0.835 0.345 Ref. Ref. 0.270 7.125 0.784 0.199 Ref. Ref. 0.207 5.312 0.756 0.137 Ref. Ref.
$\omega=0.2$ 0.301 10.100 0.996 0.598 0.151 6.994 0.992 0.385 0.145 5.249 0.989 0.274
$\omega=0.5$ 0.110 10.431 0.924 0.376 0.235 7.113 0.914 0.243 0.195 5.304 0.896 0.173
$\omega=1.0$ 0.025 10.536 0.800 0.266 0.285 7.154 0.785 0.172 0.213 5.322 0.757 0.122
$\omega=1.5$ 0.022 10.574 0.712 0.217 0.289 7.167 0.692 0.140 0.218 5.328 0.668 0.100
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
Table A.3: Summary of estimated causal effects under the scenario (c): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
@@
Confounders Logit 0.269 5.566 0.939 0.211 3.405 1.001 0.263 3.471 0.944 0.134 1.283 0.999 0.268 2.439 0.949 0.095 1.117 0.998
CBPS 0.295 5.584 0.939 0.212 3.734 1.004 0.271 3.480 0.945 0.135 1.322 1.002 0.267 2.446 0.950 0.095 1.113 1.001
RCAL CV 1.220 5.400 0.945 0.209 15.443 0.971 0.193 3.429 0.944 0.132 0.941 0.987 0.029 2.414 0.948 0.093 0.121 0.988
$\lambda=0.01$ 0.468 5.452 0.943 0.209 0.393 3.406 0.946 0.132 0.402 2.393 0.949 0.093
$\lambda=0.05$ 3.007 5.272 0.949 0.208 3.022 3.331 0.950 0.131 3.215 2.371 0.948 0.093
$\lambda=0.10$ 4.738 5.265 0.950 0.207 4.663 3.365 0.947 0.131 4.875 2.437 0.940 0.092
BRCAL PCIC 0.079 5.559 0.959 0.251 Ref. Ref. 0.205 3.473 0.960 0.146 Ref. Ref. 0.240 2.444 0.965 0.105 Ref. Ref.
$\omega=0.2$ 0.927 5.436 1.000 0.619 0.219 3.441 1.000 0.392 0.002 2.432 1.000 0.277
$\omega=0.5$ 0.220 5.513 1.000 0.392 0.072 3.464 1.000 0.248 0.170 2.440 1.000 0.175
$\omega=1.0$ 0.025 5.549 0.985 0.277 0.175 3.471 0.989 0.175 0.220 2.443 0.988 0.124
$\omega=1.5$ 0.131 5.560 0.957 0.226 0.207 3.474 0.959 0.143 0.240 2.444 0.960 0.101
Predictors Logit 1.433 6.321 0.930 0.226 1.222 1.021 1.695 3.802 0.948 0.146 1.141 1.012 1.522 2.712 0.942 0.103 1.110 1.006
CBPS 1.086 6.244 0.927 0.225 0.926 1.009 1.439 3.771 0.946 0.144 0.969 1.004 1.350 2.701 0.942 0.103 0.985 1.001
RCAL CV 2.698 5.855 0.925 0.210 2.300 0.946 2.409 3.749 0.927 0.132 1.622 0.998 2.088 2.712 0.916 0.093 1.523 1.006
$\lambda=0.01$ 2.503 6.361 0.906 0.209 2.333 3.702 0.930 0.132 2.005 2.624 0.924 0.094
$\lambda=0.05$ 3.256 5.464 0.942 0.210 3.588 3.351 0.953 0.132 3.464 2.429 0.944 0.093
$\lambda=0.10$ 4.537 5.325 0.948 0.208 4.894 3.336 0.953 0.131 4.781 2.451 0.940 0.092
BRCAL PCIC 1.173 6.190 0.932 0.228 Ref. Ref. 1.485 3.756 0.947 0.144 Ref. Ref. 1.371 2.697 0.940 0.103 Ref. Ref.
$\omega=0.2$ 1.829 5.878 1.000 0.621 1.779 3.674 1.000 0.392 1.514 2.666 1.000 0.277
$\omega=0.5$ 1.386 6.083 0.999 0.393 1.568 3.730 0.999 0.248 1.399 2.688 0.999 0.175
$\omega=1.0$ 1.230 6.164 0.975 0.278 1.504 3.750 0.982 0.175 1.380 2.693 0.978 0.124
$\omega=1.5$ 1.167 6.191 0.931 0.227 1.484 3.757 0.947 0.143 1.370 2.697 0.938 0.101
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[16]; RCAL: Calibrating method proposed by Tan[31]; BRCAL: Proposed method.
@@
Table A.4: Summary of estimated causal effects under the scenario (d): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each situation (in “Covariate” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Covariate Method
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Bias
(×$(\times 10^{-2})$
RMSE
(×$(\times 10^{-2})$
95%CP AvL BR RR
Confounders Logit 0.781 5.684 0.936 0.215 1.402 1.001 0.576 3.540 0.948 0.137 1.291 0.999 0.565 2.470 0.949 0.097 1.247 0.997
CBPS 0.716 5.711 0.937 0.216 1.285 1.006 0.493 3.553 0.949 0.138 1.105 1.002 0.472 2.479 0.949 0.098 1.042 1.000
RCAL CV 0.615 5.449 0.942 0.208 1.104 0.959 0.273 3.470 0.944 0.132 0.612 0.979 0.476 2.432 0.943 0.093 1.051 0.981
$\lambda=0.01$ 0.135 5.506 0.939 0.209 0.118 3.441 0.946 0.132 0.181 2.400 0.947 0.093
$\lambda=0.05$ 2.363 5.269 0.949 0.208 2.450 3.319 0.951 0.131 2.294 2.332 0.951 0.092
$\lambda=0.10$ 4.909 5.241 0.949 0.206 5.449 3.375 0.946 0.130 5.517 2.439 0.940 0.092
BRCAL PCIC 0.557 5.679 0.952 0.235 Ref. Ref. 0.446 3.545 0.961 0.145 Ref. Ref. 0.453 2.478 0.963 0.104 Ref. Ref.
$\omega=0.2$ 0.524 5.522 1.000 0.619 0.047 3.503 1.000 0.392 0.284 2.465 1.000 0.277
$\omega=0.5$ 0.249 5.624 1.000 0.392 0.336 3.533 0.999 0.248 0.418 2.473 0.999 0.175
$\omega=1.0$ 0.497 5.665 0.986 0.277 0.421 3.543 0.987 0.175 0.443 2.477 0.986 0.124
$\omega=1.5$ 0.576 5.679 0.952 0.227 0.447 3.546 0.959 0.143 0.453 2.478 0.958 0.101
Predictors Logit 2.977 6.551 0.927 0.233 1.119 1.019 2.961 3.989 0.943 0.150 1.085 1.010 2.796 2.803 0.944 0.107 1.068 1.004
CBPS 2.617 6.493 0.925 0.232 0.984 1.010 2.709 3.965 0.940 0.149 0.992 1.004 2.603 2.796 0.945 0.107 0.995 1.001
RCAL CV 3.792 6.057 0.921 0.210 1.426 0.942 3.770 4.038 0.901 0.132 1.381 1.023 3.634 2.900 0.890 0.093 1.389 1.039
$\lambda=0.01$ 4.470 7.089 0.879 0.209 3.752 4.012 0.901 0.132 3.353 2.783 0.904 0.094
$\lambda=0.05$ 3.510 5.532 0.943 0.210 3.288 3.427 0.944 0.132 3.200 2.436 0.945 0.093
$\lambda=0.10$ 5.066 5.278 0.952 0.207 5.194 3.365 0.944 0.130 5.460 2.464 0.937 0.092
BRCAL PCIC 2.660 6.430 0.924 0.228 Ref. Ref. 2.730 3.948 0.931 0.144 Ref. Ref. 2.617 2.792 0.932 0.102 Ref. Ref.
$\omega=0.2$ 3.060 6.054 1.000 0.621 2.890 3.857 1.000 0.392 2.705 2.761 1.000 0.277
$\omega=0.5$ 2.754 6.305 0.998 0.393 2.779 3.918 0.998 0.248 2.635 2.780 0.999 0.175
$\omega=1.0$ 2.678 6.395 0.970 0.278 2.734 3.941 0.973 0.175 2.622 2.789 0.975 0.124
$\omega=1.5$ 2.660 6.431 0.923 0.227 2.730 3.948 0.930 0.143 2.616 2.792 0.930 0.101

### v3

@@ B.2 Additional Results for Simulation Experiments
(c) Logit 1.897 6.275 0.937 0.226 1.181 1.025 1.319 3.800 0.947 0.145 1.192 1.009 1.240 2.693 0.947 0.103 1.140 1.005
CBPS 1.522 6.178 0.934 0.225 0.948 1.009 1.074 3.780 0.945 0.144 0.970 1.004 1.066 2.683 0.945 0.103 0.981 1.002
RCAL CV 2.893 5.809 0.933 0.210 1.802 0.949 1.974 3.718 0.926 0.132 1.783 0.987 1.775 2.688 0.919 0.093 1.633 1.004
$\lambda=0.01$ 2.709 6.238 0.918 0.209 1.921 3.659 0.931 0.132 1.696 2.592 0.929 0.094
$\lambda=0.05$ 3.350 5.361 0.947 0.209 3.239 3.322 0.954 0.132 3.242 2.380 0.951 0.093
$\lambda=0.10$ 4.518 5.263 0.950 0.207 4.492 3.313 0.953 0.131 4.516 2.400 0.945 0.092
BLPS 1.680 6.542 0.973 0.284 1.219 3.854 0.981 0.177 1.188 2.710 0.980 0.124
BRCAL PCIC 1.605 6.122 0.939 0.228 Ref. Ref. 1.107 3.766 0.944 0.144 Ref. Ref. 1.087 2.679 0.942 0.102 Ref. Ref.
$\omega=0.2$ 2.097 5.815 1.000 0.621 1.379 3.683 1.000 0.392 1.211 2.649 1.000 0.277
$\omega=0.5$ 1.757 6.016 0.999 0.393 1.185 3.737 0.999 0.248 1.118 2.670 0.999 0.175
$\omega=1.0$ 1.652 6.095 0.977 0.278 1.135 3.758 0.981 0.175 1.092 2.676 0.980 0.124
$\omega=1.5$ 1.606 6.123 0.939 0.227 1.110 3.766 0.944 0.143 1.088 2.679 0.940 0.101
(d) Logit 2.422 6.435 0.932 0.232 1.151 1.020 2.901 3.980 0.943 0.150 1.085 1.009 2.493 2.778 0.945 0.107 1.076 1.004
CBPS 2.047 6.365 0.932 0.232 0.973 1.009 2.653 3.961 0.941 0.150 0.993 1.004 2.299 2.773 0.944 0.107 0.992 1.002
RCAL CV 3.249 5.985 0.927 0.210 1.544 0.949 3.719 4.031 0.902 0.132 1.392 1.022 3.421 2.882 0.895 0.093 1.477 1.041
$\lambda=0.01$ 3.945 6.914 0.888 0.210 3.618 3.980 0.905 0.132 3.124 2.754 0.912 0.094
$\lambda=0.05$ 3.039 5.423 0.948 0.210 3.471 3.409 0.946 0.132 3.109 2.406 0.945 0.093
$\lambda=0.10$ 4.553 5.227 0.952 0.207 5.460 3.373 0.944 0.130 5.304 2.441 0.939 0.092
BLPS 2.210 6.757 0.969 0.285 2.823 4.044 0.972 0.177 2.447 2.799 0.973 0.125
BRCAL PCIC 2.104 6.308 0.932 0.227 Ref. Ref. 2.672 3.944 0.930 0.144 Ref. Ref. 2.317 2.768 0.934 0.102 Ref. Ref.
$\omega=0.2$ 2.507 5.970 1.000 0.621 2.842 3.851 1.000 0.392 2.404 2.736 1.000 0.277
$\omega=0.5$ 2.216 6.196 0.998 0.393 2.715 3.915 0.998 0.248 2.343 2.758 0.998 0.175
$\omega=1.0$ 2.130 6.279 0.973 0.278 2.685 3.937 0.974 0.175 2.319 2.764 0.974 0.124
$\omega=1.5$ 2.102 6.309 0.931 0.227 2.672 3.944 0.929 0.143 2.314 2.768 0.933 0.101
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[4]; RCAL: Calibrating method proposed by Tan[10]; BLPS: Bayesian method proposed by Saarela et al.[Sa2015]; BRCAL: Proposed method;
(c): “Moderate complexity and good overlap,” (d): “Moderate complexity and poor overlap.”
Table B.3: Summary of estimated causal effects under the scenario (e) and (f) (“Predictor” situation): Bias, root mean squared error (RMSE), 95% coverage probability (95%CP), average length of CI (AvL), and bias and RMSE ratio (BR/RR) compared to the bias/RMSE of our proposed method of the IPW estimator in 10,000 iterations were summarized for each scenario (in “Scenario” column) and method (in the “Method” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL: PCIC was used for the PCIC to select the learning late $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Scenario Method
@@
(e) Logit 3.150 8.186 0.910 0.266 1.234 1.044 4.181 5.107 0.935 0.181 1.185 1.030 4.084 3.555 0.942 0.132 1.209 1.018
CBPS 2.509 7.922 0.914 0.278 0.983 1.010 3.515 4.981 0.925 0.177 0.996 1.005 3.369 3.502 0.936 0.129 0.998 1.002
RCAL CV 3.367 6.785 0.886 0.209 1.319 0.865 5.009 5.329 0.827 0.132 1.419 1.075 4.810 3.915 0.788 0.093 1.424 1.121
$\lambda=0.01$ 10.530 14.410 0.728 3.457 7.313 6.792 0.743 0.133 5.235 4.102 0.766 0.093
$\lambda=0.05$ 2.902 6.812 0.886 0.210 2.740 3.937 0.912 0.132 2.412 2.722 0.916 0.094
$\lambda=0.10$ 3.353 5.514 0.938 0.207 3.984 3.452 0.940 0.130 3.997 2.454 0.940 0.092
BLPS 3.167 8.985 0.924 0.304 4.235 5.321 0.925 0.182 4.116 3.629 0.924 0.127
BRCAL PCIC 2.553 7.840 0.862 0.235 Ref. Ref. 3.529 4.958 0.861 0.147 Ref. Ref. 3.377 3.494 0.864 0.104 Ref. Ref.
$\omega=0.2$ 2.798 7.397 1.000 0.620 3.565 4.824 1.000 0.392 3.402 3.438 1.000 0.277
$\omega=0.5$ 2.639 7.710 0.988 0.391 3.541 4.917 0.988 0.248 3.387 3.478 0.987 0.175
$\omega=1.0$ 2.576 7.817 0.921 0.276 3.526 4.949 0.926 0.175 3.381 3.489 0.925 0.124
$\omega=1.5$ 2.548 7.851 0.852 0.226 3.527 4.959 0.854 0.143 3.378 3.494 0.856 0.101
(f) Logit 7.391 8.060 0.920 0.268 1.279 1.021 8.451 5.041 0.932 0.179 1.286 1.013 8.031 3.595 0.931 0.129 1.323 1.013
CBPS 5.781 7.969 0.929 0.291 1.000 1.010 6.563 4.998 0.933 0.184 0.998 1.005 6.066 3.558 0.938 0.133 1.000 1.002
RCAL CV 6.878 7.447 0.857 0.208 1.190 0.944 11.817 6.799 NA NA 1.797 1.366 12.872 5.486 2.121 1.545
$\lambda=0.01$ 19.805 19.236 0.721 292.898 17.018 9.565 0.640 0.161 14.547 6.070 0.626 0.100
$\lambda=0.05$ 8.547 8.995 0.811 0.271 7.219 4.784 0.836 0.132 6.261 3.312 0.841 0.093
$\lambda=0.10$ 4.965 5.891 0.918 0.208 5.866 3.702 0.922 0.131 5.656 2.681 0.912 0.093
BLPS 7.685 8.790 0.931 0.304 8.588 5.214 0.924 0.182 8.085 3.651 0.921 0.127
BRCAL PCIC 5.780 7.892 0.856 0.230 Ref. Ref. 6.574 4.976 0.858 0.146 Ref. Ref. 6.068 3.550 0.856 0.104 Ref. Ref.
$\omega=0.2$ 5.758 7.453 1.000 0.619 6.604 4.849 1.000 0.391 6.089 3.505 1.000 0.276
$\omega=0.5$ 5.777 7.756 0.986 0.391 6.588 4.937 0.986 0.247 6.075 3.537 0.987 0.175
$\omega=1.0$ 5.786 7.862 0.920 0.276 6.583 4.968 0.921 0.175 6.073 3.547 0.918 0.123
$\omega=1.5$ 5.783 7.894 0.850 0.225 6.576 4.976 0.850 0.143 6.071 3.550 0.844 0.101
@@
Logit: ordinary logistic regression; CBPS: Covariate balancing propensity score[4]; RCAL: Calibrating method proposed by Tan[10]; BLPS: Bayesian method proposed by Saarela et al.[Sa2015]; BRCAL: Proposed method;
(e): “Large complexity and good overlap,” (f): “Large complexity and poor overlap.”
@@
Figure B.1: Summary of estimated causal effects: Bias, root mean squared error (RMSE), and 95% coverage probability (95%CP) in 10,000 iterations were summarized under “Predictor” situation. The left panel: $n=200$; the center panel: $n=500$; The right panel: $n=1000$. RCAL was used for the CV described in Section 4.2. BRCAL was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
(a): “Small complexity and good overlap,” (b): “Small complexity and poor overlap,” (c): “Moderate complexity and good overlap,” (d): “Moderate complexity and poor overlap,” (e): “Large complexity and good overlap,” (f): “Large complexity and poor overlap.”
B.3 Covariate balance diagnosis
The results of the covariate balancing diagnostics using the standardized mean differences (SMDs) are presented in Table B.4. First, it is evident that confounding bias exists in this simulation dataset. In logistic regression, simple true propensity score settings, such as (a) and (b), achieve good covariate balance. However, as the model complexity increases ((e) and (f)), this balance is no longer maintained. As expected, the CBPS ensures exact covariate balance across all settings. In contrast, the RCAL does not necessarily achieve as strong covariate balance as other methods. In some cases, the absolute value of the SMD exceeds 0.1, which may be undesirable in causal inference contexts.
Our proposed method does not achieve exact balance like the CBPS, but it maintains relatively good covariate balance. Combined with the simulation results in the main manuscript, these findings suggest that our proposed method performs similarly to the CBPS and outperforms the RCAL in terms of covariate balance. From a causal inference perspective, our proposed method follows the concept of a regularized approach by allowing some uncertainty in covariate balancing.
Table B.4: Summary of the standardized mean difference (SMD) under the scenario (a)(f) (“Confounder” situation): SMD of the IPW estimator in 10,000 iterations were summarized for each scenario (in “Scenario” column). RCAL: CV was used for the CV described in Section 4.2. BRCAL was used for the PCIC to select the learning rate $\omega$ from candidates: $(0.2,0.5,1.0,1.5)$ (see Section 4.2).
Scenario Variables SMD
Before Logit CBPS RCAL BLPS BRCAL Before Logit CBPS RCAL BLPS BRCAL Before Logit CBPS RCAL BLPS BRCAL
(a) $X_{1}$ 0.271 0.001 0.000 0.080 -0.012 0.014 0.268 0.000 0.000 0.018 -0.005 0.005 0.268 0.000 0.000 0.007 -0.002 0.002
$X_{2}$0.318 0.000 0.000 -0.087 0.015 -0.015 -0.316 0.000 0.000 -0.018 0.006 -0.005 -0.315 0.000 0.000 -0.007 0.003 -0.002
$X_{3}$ 0.251 0.001 0.000 0.076 -0.011 0.013 0.251 0.000 0.000 0.018 -0.004 0.005 0.252 0.000 0.000 0.006 -0.002 0.002
$X_{4}$0.169 0.000 0.000 -0.057 0.008 -0.010 -0.171 0.000 0.000 -0.015 0.003 -0.004 -0.169 0.000 0.000 -0.006 0.001 -0.002
(b) $X_{1}$ 0.433 0.005 0.000 0.085 -0.019 0.008 0.433 0.004 0.000 0.029 -0.006 0.003 0.432 0.004 0.000 0.015 -0.001 0.002
$X_{2}$0.520 -0.003 $0$0.087 0.026 -0.008 -0.518 -0.002 $0$0.029 0.010 -0.003 -0.518 -0.002 $0$0.015 0.004 -0.002
$X_{3}$ 0.407 0.005 0.000 0.083 -0.018 0.008 0.406 0.003 0.000 0.028 -0.006 0.003 0.405 0.003 0.000 0.014 -0.001 0.002
$X_{4}$0.269 -0.004 $0$0.073 0.011 -0.007 -0.269 -0.003 $0$0.026 0.003 -0.003 -0.270 -0.003 $0$0.013 $0$0.002
(c) $X_{1}$ 0.289 0.015 0.000 0.121 0.002 0.014 0.286 0.015 0.000 0.072 0.010 0.005 0.285 0.015 0.000 0.061 0.012 0.002
$X_{2}$0.027 -0.002 $0$0.013 $0$0.002 -0.030 -0.001 $0$0.008 -0.001 -0.001 -0.031 -0.001 $0$0.007 -0.001 -0.001
$X_{3}$ 0.345 0.012 0.000 0.121 -0.003 0.014 0.344 0.012 0.000 0.062 0.006 0.005 0.344 0.012 0.000 0.050 0.009 0.002
$X_{4}$0.272 -0.002 $0$0.089 0.011 -0.013 -0.271 -0.002 $0$0.035 0.003 -0.005 -0.270 -0.002 $0$0.022 0.001 -0.002
(d) $X_{1}$ 0.292 0.019 0.000 0.138 0.005 0.011 0.293 0.019 0.000 0.099 0.014 0.004 0.292 0.019 0.000 0.089 0.016 0.002
$X_{2}$0.114 -0.003 $0$0.046 0.003 -0.006 -0.113 -0.003 $0$0.023 $0$0.003 -0.116 -0.003 $0$0.017 -0.001 -0.002
$X_{3}$ 0.392 0.013 0.000 0.128 -0.005 0.012 0.391 0.013 0.000 0.072 0.006 0.004 0.393 0.013 0.000 0.059 0.010 0.002
$X_{4}$0.316 -0.002 $0$0.088 0.013 -0.011 -0.314 -0.002 $0$0.031 0.004 -0.004 -0.312 -0.002 $0$0.017 0.001 -0.002
(e) $X_{1}$ 0.237 0.003 0.000 0.112 -0.010 0.007 0.235 0.002 0.000 0.065 -0.003 0.003 0.235 0.002 0.000 0.048 -0.001 0.002
$X_{2}$0.643 -0.015 $0$0.061 0.020 -0.008 -0.639 -0.014 0.000 0.029 $0$0.003 -0.638 -0.013 0.000 0.054 -0.007 -0.002
$X_{3}$ 0.273 0.002 0.000 0.105 -0.013 0.007 0.275 0.001 0.000 0.048 -0.005 0.003 0.274 0.001 0.000 0.029 -0.003 0.002
$X_{4}$0.453 -0.019 $0$0.044 0.005 -0.008 -0.449 -0.019 0.000 0.041 -0.009 -0.003 -0.448 -0.018 0.000 0.065 -0.014 -0.002
(f) $X_{1}$ 0.146 0.007 0.000 0.116 -0.002 0.004 0.143 0.006 0.000 0.098 0.002 0.002 0.142 0.005 0.000 0.085 0.004 0.001
$X_{2}$0.736 -0.052 $0$0.055 -0.011 -0.008 -0.737 -0.050 0.000 0.089 -0.034 -0.003 -0.736 -0.049 0.000 0.144 -0.041 -0.002
$X_{3}$ 0.225 0.001 0.000 0.129 -0.012 0.006 0.225 0.000 0.000 0.074 -0.006 0.003 0.224 -0.001 0.000 0.043 -0.004 0.001
$X_{4}$0.576 -0.061 $0$0.039 -0.030 -0.007 -0.572 -0.058 0.000 0.100 -0.046 -0.003 -0.572 -0.058 0.000 0.153 -0.052 -0.002